For a given gas the coefficient of viscosity is 1.9 × 105 N s m2 and the diffusion coefficient is 1.2 × 105 m 2 s 1 . Calculate the density and mean free path at average molecular velocity of 380 m s1 . 


1
Expert's answer
2021-04-08T07:36:12-0400



First we recall all the terms and their calculation formula to drieve our eqn..




We know, that,




1)Mean free path "=\boxed{\lambda={RT \over \sqrt 2 \pi d^2 N_A P}}"

2)And also given by


"\boxed{\lambda={{\mu \over P}{ \sqrt {\pi k T \over 2 m}}}}"


"\u200b"


 "\mu=1.9\times 19^5" N s m2

p = pressure

T = temperature

m = molecular mass

 

  






And 




3)Average velocity is given by

"V=\sqrt{kT\over m \pi}"


 

 

 


4)And Diffusion coffecient is calculated by


"D={kT\over 6\pi \mu R_o}"



Given:

"\bigstar" coefficient of viscosity is 1.9 × 10-5 N s m2


and


"\bigstar" diffusion coefficient is 1.2 × 10-5 m2 s-1


"\bigstar" average molecular velocity is 380 ms-1



now after manipulations we get,



"{6D\over K}={T\over \pi \mu}={\pi d^2 N_A\over R}"

on comparing we get density,

"{\sqrt {6DR\over K N_A \pi}}={ d }"

"\boxed{d=1.55\times10^{-6}}"


and we also get mean free path ..

"\boxed{\lambda =6.623 \times 10^{-3}}"